OpenStudy (anonymous):

how do you find a particular solution to (15/4)y''-4y'+y=3xe^2x

OpenStudy (turingtest):

undetermined coefficients is not too painful

OpenStudy (anonymous):

i solved it but got it wrong

OpenStudy (turingtest):

our guess will be\[Y_p=(Ax+B)e^{2x}\]\[Y'_p=(2Ax+A+B)e^{2x}\]\[Y''_p=(4Ax+4A+B)e^{2x}\]

OpenStudy (anonymous):

can i send you my work so you can see where I went wrong?

OpenStudy (turingtest):

ok, but i'm gonna keep going in the meantime

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

TuringTest - your choice of a particular solution is correct, but I think your expressions for the first and second derivatives are wrong.

OpenStudy (turingtest):

yes I know I messed up on B I'm fixing it...

OpenStudy (asnaseer):

ok - rilayian I'll leave you in the capable hands of TuringTest :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I attached my work above and I've looked over it twice and I'm still not seeing where i went wrong'

OpenStudy (turingtest):

@asnaseer thanks :D @rilayian that's kinda difficult to read...

OpenStudy (anonymous):

I'll send it again

OpenStudy (anonymous):

I'm going to send it on 3 parts

OpenStudy (turingtest):

you never plugged it into the differential equation!

OpenStudy (turingtest):

you have to take your formulas for y, y', and y'' and put them into your problem, then you can solve the system to find A and B

OpenStudy (anonymous):

i just realized that lol but even when i did I still got it wrong

OpenStudy (turingtest):

your work is fine besides that as far as I can see

OpenStudy (anonymous):

i ended up with (33/64)e^2x +(3/8)xe^2x

OpenStudy (anonymous):

and it's still wrong

OpenStudy (turingtest):

33/64 should be negative

OpenStudy (anonymous):

omg thank you soo much!!

OpenStudy (turingtest):

should have been\[Y'_p=(2Ax+A+2B)e^{2x}\]\[Y''_p=(4Ax+4A+4B)e^{2x}\]plug it into the equation:\[(4Ax+4A+4B)e^{2x}-\frac{16}{15}(2Ax+A+2B)e^{2x}+\frac{4}{15}(Ax+B)e^{2x}\]and you get for the Ax part\[4Ax-\frac{32}{15}Ax+\frac4{15}Ax=\frac45x\]\[60A-32A+4=12\to A=\frac38\]this is what I had so far, you can finish it out I think if you want

OpenStudy (turingtest):

my A is your B

OpenStudy (turingtest):

and I already have the coefficient of y'' set to 1, so my way is different

OpenStudy (anonymous):

thanks but once I changed it to negative it ended up being correct! thanks soo much

OpenStudy (turingtest):

welcome!

OpenStudy (anonymous):

:)